# Free Doubly-Infinitary Distributive Categories are Cartesian Closed

### v6

@@ 2. The Free Doubly-Infinitary Distributive Category
$\left(\left[\langle\prescript{k}{}{C}\left(j,i\right)\mid i\in\prescript{k}{}{I}_{j}\rangle\mid j\in\prescript{k}{}{J}\right]\right)_{k\in K},$
@@
$\coprod_{k\in K}\left[\langle\prescript{k}{}{C}\left(j,i\right)\mid i\in\prescript{k}{}{I}_{j}\rangle\mid j\in\prescript{k}{}{J}\right]\cong\left[\langle\prescript{t}{}{C}\left(j,i\right)\mid i\in\prescript{t}{}{I}_{j}\rangle\mid(t,j)\in\displaystyle\coprod_{k\in K}\left(\left\{k\right\}\times\prescript{k}{}{J}\right)\right].$
@@
Since $\mathbf{Dist}\left(\mathbb{C}\right)$ is of the form $\mathbf{Fam}(\mathbb{D})$ for $\mathbb{D}=\mathbf{Fam}(\mathbb{C}^{op})^{op}$, and coproducts in $\mathbf{Fam}(\mathbb{D})$ are freely generated, coproducts exist by construction. ∎
@@
$\left(\left[\langle\prescript{k}{}{C}\left(j,i\right)\mid i\in\prescript{k}{}{I}_{j}\rangle\mid j\in\prescript{k}{}{J}\right]\right)_{k\in K},$
@@
$\prod_{k\in K}\left[\langle\prescript{k}{}{C}\left(j,i\right)\mid i\in\prescript{k}{}{I}_{j}\rangle\mid j\in\prescript{k}{}{J}\right]\cong\left[\langle\prescript{t}{}{C}\left(j_{t},i\right)\mid\left(t,i\right)\in\coprod_{k\in K}\left(\left\{k\right\}\times\prescript{k}{}{I}_{j}\right)\rangle\mid j\in\prod_{k\in K}\prescript{k}{}{J}\right]$
@@
$\prod_{k\in K}\left[\prescript{k}{}{D}(j)\mid j\in\prescript{k}{}{J}\right]\cong\left[\prod_{t\in K}\prescript{k}{}{D}(j_{t})\mid j\in\prod_{k\in K}\prescript{k}{}{J}\right].$
@@
$\prod_{t\in K}\langle\prescript{t}{}{C}\left(j_{t},i\right)\mid i\in\prescript{t}{}{I}_{j}\rangle\cong\langle\prescript{t}{}{C}\left(j_{t},i\right)\mid\left(t,i\right)\in\coprod_{k\in K}\left(\left\{k\right\}\times\prescript{k}{}{I}\right)\rangle.$
Exponentials in $\mathbf{Dist}(\mathbb{C})$
$\mathbf{Dist}\left(\mathbb{C}\right)$ is a cartesian closed category. Actually, the construction of the exponential is reminiscent of a Dialectica-like interpretation [18, 26, 33].
@@
The category $\mathbf{Dist}{\mathbb{C}}$ is cartesian closed. Exponentials are given by the following formula. Given
$A=\left[\langle C(j,i)\mid i\in I_{j}\rangle\mid j\in J\right],\qquad B=\left[\langle C^{\prime}(j^{\prime},i^{\prime})\mid i^{\prime}\in I^{\prime}_{j^{\prime}}\rangle\mid j^{\prime}\in J^{\prime}\right],$
@@
(2.6) $\coprod_{j^{\prime}\in J^{\prime}}\prod_{i^{\prime}\in I^{\prime}_{j^{\prime}}}\prod_{j\in J}\left[\langle F^{(j^{\prime},i^{\prime})}(t,r)\mid t\in\left\{\ast\right\}\rangle\mid r\in\left(\left\{\ast\right\}\sqcup\mathbf{Dist}\left(\mathbb{C}\right)\left(\underline{\langle C(j,i)\mid i\in I_{j}\rangle},\overline{C^{\prime}(j^{\prime},i^{\prime})}\right)\right)\right]$
where: (1) we denote by $\underline{\langle C(j,i)\mid i\in I_{j}\rangle}$ the image of the object $\langle C(j,i)\mid i\in I_{j}\rangle$ by the inclusion $\mathbf{Fam}\left(\mathbb{C}^{\mathrm{op}}\right)^{\mathrm{op}}\to\mathbf{Dist}\left(\mathbb{C}\right)$; (2) we denote by $\overline{C^{\prime}(j^{\prime},i^{\prime})}$ the image of $C^{\prime}(j^{\prime},i^{\prime})$ by the inclusion $\mathbb{C}\to\mathbf{Dist}\left(\mathbb{C}\right)$; (3) $F^{(j^{\prime},i^{\prime})}(\ast,\ast)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}C^{\prime}(j^{\prime},i^{\prime})$ and, in case $r\neq\ast$, $F(\ast,r)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathbb{1}$ where we denote by $\mathbb{1}$ the terminal object in $\mathbf{Dist}\left(\mathbb{C}\right)$.
@@
The proof consists of establishing a natural isomorphism of hom-sets defining the exponential adjunction. This proceeds by rewriting the hom-sets of $\mathbf{Dist}(\mathbb{C})$, using explicit descriptions of coproducts and products in $\mathbf{Fam}(\mathbb{C})$ and $\mathbf{Fam}(\mathbb{C}^{op})^{op}$ as given above. The argument follows standard rearrangements of products and coproducts indexed over finite sums, and repeatedly applies distributivity of dependent sums over products (in a purely set-theoretic sense). Full details follow the same pattern as the calculation provided in the dependent type version. We refer the reader to [26, Section 4.2] for technical details in a more general setting, otherwise, for an inductive approach, we refer the reader to Remark 2.6 and the proof of [32, Theorem 3.1] – which is entirely analogous. ∎
@@
(2.7) $\displaystyle\coprod_{j^{\prime}\in J^{\prime}}\prod_{i^{\prime}\in I^{\prime}_{j^{\prime}}}\left[\langle F^{(j^{\prime},i^{\prime})}(j,r_{j})\mid j\in J\rangle\mid r\in\prod_{j\in J}\left(\left\{\ast\right\}\sqcup\mathbf{Dist}\left(\mathbb{C}\right)\left(\underline{\langle C(j,i)\mid i\in I_{j}\rangle},\overline{C^{\prime}(j^{\prime},i^{\prime})}\right)\right)\right]$
@@
(2.8) $\displaystyle\coprod_{j^{\prime}\in J^{\prime}}\left[\langle F^{(j^{\prime},t)}(j,r_{(t,j)})\mid(t,j)\in\prod_{i^{\prime}\in I^{\prime}_{j^{\prime}}}\left(\left\{i^{\prime}\right\}\times J\right)\rangle\mid r\in\prod_{(i^{\prime},j)\in I^{\prime}_{j^{\prime}}\times J}\left(\left\{\ast\right\}\sqcup\mathbf{Dist}\left(\mathbb{C}\right)\left(\underline{\langle C(j,i)\mid i\in I_{j}\rangle},\overline{C^{\prime}(j^{\prime},i^{\prime})}\right)\right)\right]$
@@
(2.9) $\displaystyle\left[\langle F^{(k,t)}(j,r_{(t,j)})\mid(t,j)\in\prod_{i^{\prime}\in I^{\prime}_{k}}\left(\left\{i^{\prime}\right\}\times J\right)\rangle\mid\left(k,r\right)\in\coprod_{j^{\prime}\in J^{\prime}}\left(\left\{j^{\prime}\right\}\times\prod_{(i^{\prime},j)\in I^{\prime}_{j^{\prime}}\times J}\left(\left\{\ast\right\}\sqcup\mathbf{Dist}\left(\mathbb{C}\right)\left(\underline{\langle C(j,i)\mid i\in I_{j}\rangle},\overline{C^{\prime}(j^{\prime},i^{\prime})}\right)\right)\right)\right]$
@@
    $A\Rightarrow B=\left(\coprod_{i\in\mathbf{Dist}{\mathbb{C}}(A,B)}\mathbb{1}\right)\sqcup B.$
@@
The above suffices to define exponentials by structural induction, since both functor $A\Rightarrow-:\mathbf{Dist}(\mathbb{C})\to\mathbf{Dist}(\mathbb{C})$ and $-\Rightarrow A:\mathbf{Dist}(\mathbb{C})^{\mathrm{op}}\to\mathbf{Dist}(\mathbb{C})$ preserve products.
@@
The argument of Theorem 2.4 also shows that in the free (non-doubly) infinitary distributive category $\mathbf{Fam}(\mathbf{FinFam}(\mathbb{C}^{op})^{op})$, exponentials $A\Rightarrow B$ exist whenever $A$ belongs to the image of $\mathbf{FinFam}(\mathbf{FinFam}(\mathbb{C}^{op})^{op})\to\mathbf{Fam}(\mathbf{FinFam}(\mathbb{C}^{op})^{op})$.
@@ 3. The Pseudomonad Structure of $\mathbf{Dist}$
$\langle\left[C(j,i)\mid i\in I_{j}\right]\mid j\in J\rangle.$
@@
$\left[\langle C\left(j,r_{j}\right)\mid j\in J\rangle\mid r\in\prod_{t\in J}I_{t}\right].$
@@
By construction, the collection of morphisms $\lambda_{\mathbb{C}}$ is pseudonatural in $\mathbb{C}$. This pseudonaturality is a consequence of the universal properties of free completions and the naturality of the pseudomonad units on $\mathbf{Cat}$.
More precisely, following [48, Theorem 7.1] with $\mathbb{B}=\mathbf{Set}$, these data define a pseudodistributive law
@@
in the sense of [37, Definition 11.4], and this law is essentially unique [49, Remark 34, Corollary 49].
@@
$\mathbf{Dist}$-pseudoalgebras
Next, we consider the pseudoalgebras of $\mathbf{Dist}$. We start by defining:
@@
Let $\mathbb{C}$ be a category with products and coproducts. Then, for each family of objects $\left(C_{ij}\right)_{\left(j,i\right)\in J\times I_{j}}$ in $\mathbb{C}$, we have a canonical morphism
(3.1) $[\displaystyle\langle\iota_{r_{j}}\circ\pi_{j}\mid j\in J\rangle\mid r\in\Pi_{j\in J}I_{j}]:\left(\coprod_{r\in\Pi_{j\in J}I_{j}}\prod_{j\in J}C_{r_{j}j}\right)\to\left(\prod_{j\in J}\coprod_{i\in I_{j}}C_{ij}\right).$
We say that $\mathbb{C}$ is doubly-infinitary distributive (i.e., that infinite products distribute over infinite coproducts in $\mathbb{C}$) if (3.1) is invertible for any family $\left(\left(C_{ij}\right)_{i\in I_{j}}\right)_{j\in J}$.
@@
By the explicit construction of products in $\mathbf{Fam}(\mathbb{C})$ (see Section 1), this compatibility condition precisely corresponds to the invertibility of the canonical morphisms characterizing doubly-infinitary distributivity as defined in Def. 3.1.
@@
that gives a lifting (as defined in [35]) of the free completion under coproducts pseudomonad to $\mathbf{ProdCat}$.

### v7

@@ 2.2. Products and coproducts in $\mathbf{Dist}\left({\mathbb{C}}\right)$
$\left(\left[\langle\prescript{k}{}{C}\left(j,i\right)|i\in\prescript{k}{}{I}_{j}\rangle|j\in\prescript{k}{}{J}\right]\right)_{k\in K},$
@@
$\coprod_{k\in K}\left[\langle\prescript{k}{}{C}(j,i)|i\in\prescript{k}{}{I}_{j}\rangle|j\in\prescript{k}{}{J}\right]\cong\left[\langle\prescript{t}{}{C}(j,i)|i\in\prescript{t}{}{I}_{j}\rangle|(t,j)\in\coprod_{k\in K}\left(\left\{k\right\}\times\prescript{k}{}{J}\right)\right].$
@@
Since $\mathbf{Dist}(\mathbb{C})$ is of the form $\mathbf{Fam}(\mathbb{D})$, where $\mathbb{D}=\mathbf{Fam}(\mathbb{C}^{op})^{op}$, and coproducts in $\mathbf{Fam}(\mathbb{D})$ are freely adjoined (i.e., formal coproducts), it follows that coproducts exist by construction. ∎
@@
$\left(\left[\langle\prescript{k}{}{C}\left(j,i\right)|i\in\prescript{k}{}{I}_{j}\rangle|j\in\prescript{k}{}{J}\right]\right)_{k\in K},$
@@
$\prod_{k\in K}\left[\langle\prescript{k}{}{C}(j,i)|i\in\prescript{k}{}{I}_{j}\rangle|j\in\prescript{k}{}{J}\right]\cong\left[\langle\prescript{t}{}{C}(j_{t},i)|(t,i)\in\coprod_{k\in K}\left\{k\right\}\times\prescript{k}{}{I}_{j_{k}}\rangle|j\in\prod_{k\in K}\prescript{k}{}{J}\right].$
@@
$\prod_{k\in K}\left[\prescript{k}{}{D}(j)|j\in\prescript{k}{}{J}\right]\cong\left[\prod_{k\in K}\prescript{k}{}{D}(j_{k})|j\in\prod_{k\in K}\prescript{k}{}{J}\right].$
@@
$\prod_{k\in K}\langle\prescript{k}{}{C}(j_{k},i)|i\in\prescript{k}{}{I}_{j_{k}}\rangle\;\cong\;\langle\prescript{t}{}{C}(j_{t},i)|(t,i)\in\coprod_{k\in K}\left\{k\right\}\times\prescript{k}{}{I}_{j_{k}}\rangle.$
2.3. Exponentials in $\mathbf{Dist}(\mathbb{C})$
$\mathbf{Dist}\left(\mathbb{C}\right)$ is a cartesian closed category. Actually, the construction of the exponential is reminiscent of a Dialectica-like interpretation [20, 17, 51, 29, 37].
@@ 2.3. Exponentials in $\mathbf{Dist}(\mathbb{C})$
The category $\mathbf{Dist}(\mathbb{C})$ is cartesian closed. Exponentials are given by the following formula. Given
$A=\left[\langle C(j,i)|i\in I_{j}\rangle|j\in J\right],\qquad B=\left[\langle C^{\prime}(j^{\prime},i^{\prime})|i^{\prime}\in I^{\prime}_{j^{\prime}}\rangle|j^{\prime}\in J^{\prime}\right],$
@@
(2.6) $\coprod_{j^{\prime}\in J^{\prime}}\prod_{i^{\prime}\in I^{\prime}_{j^{\prime}}}\prod_{j\in J}\left[\langle F^{(j^{\prime},i^{\prime})}(t,r)|t\in\left\{\ast\right\}\rangle|r\in\left(\left\{\ast\right\}\sqcup\mathbf{Dist}\left(\mathbb{C}\right)\left(\underline{\langle C(j,i)|i\in I_{j}\rangle},\overline{C^{\prime}(j^{\prime},i^{\prime})}\right)\right)\right]$
where: (1) we denote by $\underline{\langle C(j,i)|i\in I_{j}\rangle}$ the image of the object $\langle C(j,i)|i\in I_{j}\rangle$ by the inclusion $\mathbf{Fam}\left(\mathbb{C}^{\mathrm{op}}\right)^{\mathrm{op}}\to\mathbf{Dist}\left(\mathbb{C}\right)$; (2) we denote by $\overline{C^{\prime}(j^{\prime},i^{\prime})}$ the image of $C^{\prime}(j^{\prime},i^{\prime})$ by the inclusion $\mathbb{C}\to\mathbf{Dist}\left(\mathbb{C}\right)$; (3) $F^{(j^{\prime},i^{\prime})}(\ast,\ast)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}C^{\prime}(j^{\prime},i^{\prime})$ and, in case $r\neq\ast$, $F^{(j^{\prime},i^{\prime})}(\ast,r)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathbb{1}$ where we denote by $\mathbb{1}$ the terminal object in $\mathbf{Dist}\left(\mathbb{C}\right)$.
@@
The proof proceeds by establishing a natural isomorphism of hom-sets defining the exponential adjunction. Full details follow the same pattern as the calculation provided in the dependent type version. For technical details in a more general setting, we refer the reader to [29, Section 4.2]. Alternatively, for an inductive approach, we refer the reader to Remark 2.6 and the proof of [36, Theorem 3.1], which is entirely analogous. ∎
@@
(2.7) $\displaystyle\coprod_{j^{\prime}\in J^{\prime}}\prod_{i^{\prime}\in I^{\prime}_{j^{\prime}}}\left[\langle F^{(j^{\prime},i^{\prime})}(\ast,r_{j})|j\in J\rangle\Bigg|(r_{j})_{j\in J}\in\prod_{j\in J}\left(\left\{\ast\right\}\sqcup\mathbf{Dist}\left(\mathbb{C}\right)\left(\underline{\langle C(j,i)|i\in I_{j}\rangle},\overline{C^{\prime}(j^{\prime},i^{\prime})}\right)\right)\right]$
@@
(2.8) $\displaystyle\coprod_{j^{\prime}\in J^{\prime}}\left[\langle F^{(j^{\prime},i^{\prime})}(\ast,r_{i^{\prime},j})|(i^{\prime},j)\in\coprod_{i^{\prime}\in I^{\prime}_{j^{\prime}}}J\rangle\Bigg|(r_{i^{\prime},j})_{(i^{\prime},j)\in I^{\prime}_{j^{\prime}}\times J}\in\prod_{(i^{\prime},j)\in I^{\prime}_{j^{\prime}}\times J}\left(\left\{\ast\right\}\sqcup\mathbf{Dist}\left(\mathbb{C}\right)\left(\underline{\langle C(j,i)|i\in I_{j}\rangle},\overline{C^{\prime}(j^{\prime},i^{\prime})}\right)\right)\right]$
@@
(2.9) $\displaystyle\left[\langle F^{(j^{\prime},i^{\prime})}(\ast,r_{i^{\prime},j})|(i^{\prime},j)\in\coprod_{i^{\prime}\in I^{\prime}_{j^{\prime}}}J\rangle\Bigg|(j^{\prime},(r_{i^{\prime},j})_{(i^{\prime},j)\in I^{\prime}_{j^{\prime}}\times J})\in\coprod_{j^{\prime}\in J^{\prime}}\prod_{(i^{\prime},j)\in I^{\prime}_{j^{\prime}}\times J}\left(\left\{\ast\right\}\sqcup\mathbf{Dist}\left(\mathbb{C}\right)\left(\underline{\langle C(j,i)|i\in I_{j}\rangle},\overline{C^{\prime}(j^{\prime},i^{\prime})}\right)\right)\right]$
@@
    $A\Rightarrow B=\left(\coprod_{i\in\mathbf{Dist}(\mathbb{C})(A,B)}\mathbb{1}\right)\sqcup B.$
@@
The above suffices to define exponentials by structural induction, since both functors
$(A\Rightarrow-):\mathbf{Dist}(\mathbb{C})\to\mathbf{Dist}(\mathbb{C})$  and   $(-\Rightarrow A):\mathbf{Dist}(\mathbb{C})^{\mathrm{op}}\to\mathbf{Dist}(\mathbb{C})$
preserve products.
@@
It is easy to see that the exponential formula given by Remark 2.6 can also be applied in the case of free infinitary distributive categories on categories. More precisely, in the free infinitary distributive category $\mathbf{Fam}(\mathbf{FinFam}(\mathbb{C}^{op})^{op})$, the exponential $A\Rightarrow B$ exists provided that $A$ belongs to the image of $\mathbf{FinFam}(\mathbf{FinFam}(\mathbb{C}^{op})^{op})\to\mathbf{Fam}(\mathbf{FinFam}(\mathbb{C}^{op})^{op})$.
@@ 3. The Pseudomonad Structure of $\mathbf{Dist}$
$\langle\left[C(j,i)|i\in I_{j}\right]|j\in J\rangle.$
@@
$\left[\langle C\left(j,r_{j}\right)|j\in J\rangle|r\in\prod_{t\in J}I_{t}\right].$
@@
By construction, the collection of morphisms $\lambda_{\mathbb{C}}$ is pseudonatural in $\mathbb{C}$. This pseudonaturality is a consequence of the universal properties of free completions and the pseudonaturality of the pseudomonad units on $\mathbf{Cat}$. More precisely, following [55, Theorem 7.1] with $\mathbb{B}=\mathbf{Set}$, these data define a pseudodistributive law
@@
in the sense of [41, Definition 11.4], and this law is essentially unique [56, Remark 34, Corollary 49].
@@
Doubly-infinitary distributive categories: $\mathbf{Dist}$-pseudoalgebras
We study the $\mathbf{Dist}$-pseudoalgebras, starting by describing them explicitly. We start by defining doubly-infinitary distributive categories explicitly.
@@
Let $\mathbb{C}$ be a category with products and coproducts. Then, for each family of objects $\left(C_{ji}\right)_{\left(j,i\right)\in J\times I_{j}}$ in $\mathbb{C}$, we have a canonical morphism
(3.1) $\langle\iota_{r_{j}}\circ\pi_{j}\mid j\in J\rangle\mid r\in\prod_{j\in J}I_{j}]:\left(\coprod_{r\in\prod\limits_{j\in J}I_{j}}\prod_{j\in J}C_{\,j,\,r_{j}}\right)\longrightarrow\left(\prod_{j\in J}\coprod_{i\in I_{j}}C_{\,ji}\right).$
We say that $\mathbb{C}$ is doubly-infinitary distributive (i.e., that infinite products distribute over infinite coproducts in $\mathbb{C}$) if (3.1) is invertible for any family $\left(\left(C_{ji}\right)_{i\in I_{j}}\right)_{j\in J}$.
@@
By the explicit construction of products in $\mathbf{Fam}(\mathbb{C})$, this compatibility condition precisely corresponds to the invertibility of the canonical morphisms characterizing doubly-infinitary distributivity as defined in Def. 3.1.
@@
that give a lifting (as defined in [39]) of the free coproduct completion pseudomonad on $\mathbf{ProdCat}$.
4. Free doubly-infinitary distributive categories on discrete categories
We have explored the concept of doubly-infinitary distributive categories by establishing basic results for the freely generated ones. Before turning to concrete examples, we first consider the free doubly-infinitary distributive categories on discrete categories.
In what follows, we make use of bicategorical limits (or bilimits). We refer the reader to [49, 50, 30, 33] for the relevant definitions.
4.1. Bicategorically semi-additive $2$-categories
We begin by observing that the $2$-category $\mathbf{ProdCat}$, the $2$-category of categories with products, naturally carries an enrichment over the $2$-category of symmetric monoidal categories endowed with the multilinear multicategorical structure, and that it admits products. Taken together, these facts imply that $\mathbf{ProdCat}$ is bicategorically semi-additive (see, for instance, [28] for the notion of a semi-additive category). More precisely:
Definition 4.1 (Bicategorically semi-additive $2$-categories).
A $2$-category $\mathbb{C}$ is bicategorically semi-additive if (1) it admits an enrichment over the $2$-category of symmetric monoidal categories (with the multilinear multicategorical structure); and (2) it has binary bicategorical (enriched) products as well as a bicategorical (enriched) zero object: in other words, for each pair of objects $A,B$ in $\mathbb{C}$, we have an object $A\underline{\times}B$ such that we have a pseudonatural equivalences
$\mathbb{C}\left(-,A\right)\times\mathbb{C}\left(-,B\right)\simeq\mathbb{C}\left(-,A\underline{\times}B\right)$
that is also (pointwise) monoidal. And we have an object $\top$ such that $\mathbb{C}\left(-,\top\right)\simeq\mathsf{1}$.
Definition 4.2 (Bicategorical biproduct).
Let $\mathbb{C}$ be a $2$-category with a bicategorical zero object 0, as well as binary bicategorical products $\times$ and coproducts $\sqcup$. In this setting, there is a canonical comparison morphism
(4.1) $\mathfrak{c}:A\sqcup B\to A\times B$
induced by the maps $A\to A\times B$ and $B\to A\times B$, respectively defined by $\langle{\mathrm{id}}_{A},0\rangle$ and $\langle 0,{\mathrm{id}}_{B}\rangle$, where 0 denotes the zero morphism.
Whenever $\mathfrak{c}$ is an equivalence, we say that $A\sqcup B\simeq A\times B$ is a bicategorical biproduct.
Theorem 4.3.
Any bicategorically semi-additive bicategory has bicategorical biproducts.
Proof.
The inverse (up to equivalence) of the comparison morphism (4.1) is given by
$\left(\iota_{A}\circ\pi_{A}\right)\otimes\left(\iota_{B}\circ\pi_{B}\right),$
where $\iota_{A}:A\to A\sqcup B$ and $\iota_{B}:B\to A\sqcup B$ are the coprojections of the bicategorical coproduct, and $\otimes$ denotes the monoidal product in the symmetric monoidal category $\mathbb{C}(A\times B,A\sqcup B)$. ∎
Theorem 4.4 ($\mathbf{ProdCat}$ is bicategorically semi-additive).
The $2$-category $\mathbf{ProdCat}$ is bicategorically semi-additive.
Proof.
$\mathbf{ProdCat}\to\mathbf{Cat}$
is pseudomonadic and therefore creates bicategorical products.
The pointwise product of functors induces the monoidal structure on the hom-categories of $\mathbf{ProdCat}$. ∎
As a consequence, $\mathbf{ProdCat}$ admits bicategorical biproducts. In other words:
Corollary 4.5.
The $2$-category $\mathbf{ProdCat}$ has bicategorical biproducts. More precisely, the bicategorical coproducts are given, up to equivalence, by the bicategorical products in $\mathbf{ProdCat}$.
The observations above also apply to the case of the $2$-category $\mathsf{Fin}\mathbf{ProdCat}$ of categories with finite products. That is to say:
Theorem 4.6.
The $2$-category $\mathsf{Fin}\mathbf{ProdCat}$ has bicategorical biproducts.
Proof.
The underlying category of the binary biproduct of two objects $A,B$ in $\mathsf{Fin}\mathbf{ProdCat}$ is given by the usual product of categories $A\times B$, since
$\mathsf{Fin}\mathbf{ProdCat}\to\mathbf{Cat}$
is pseudomonadic. ∎
4.2. The construction
With the observations above, we can proceed and construct freely generated doubly-infinitary distributive categories on discrete categories. More precisely:
Theorem 4.7.
The free doubly-infinitary distributive category on the discrete category with two objects is equivalent to $\mathbf{Fam}\left(\mathbf{Set}^{op}\times\mathbf{Set}^{op}\right)$.
Proof.
Indeed, we have the following equivalences
$\mathbf{Dist}\left(\mathbb{1}\sqcup_{\mathbf{Cat}}\mathbb{1}\right)$ $\simeq\mathbf{Fam}\left(\mathbf{Fam}\left(\left(\mathbb{1}\sqcup_{\mathbf{Cat}}\mathbb{1}\right)^{op}\right)^{op}\right)$
$\simeq\mathbf{Fam}\left(\mathbf{Fam}\left(\mathbb{1}^{op}\right)^{op}\sqcup_{\mathbf{ProdCat}}\mathbf{Fam}\left(\mathbb{1}^{op}\right)^{op}\right)$
$\simeq\mathbf{Fam}\left(\mathbf{Fam}\left(\mathbb{1}^{op}\right)^{op}\times\mathbf{Fam}\left(\mathbb{1}^{op}\right)^{op}\right)$
$\simeq\mathbf{Fam}\left(\mathbf{Set}^{op}\times\mathbf{Set}^{op}\right),$
where the coproduct and product signs always represent bicategorical coproducts and products in the respective $2$-categories, and we use the facts that:
  1. (1)
    the functor $op\circ\mathbf{Fam}\circ op:\mathbf{Cat}\to\mathbf{ProdCat}$ is a left biadjoint and thus preserves bicategorical coproducts,
  2. (2)
    bicategorical coproducts and products coincide in $\mathbf{ProdCat}$.
Theorem 4.7 can be further generalized to any discrete category; namely, by observing that $\mathbf{ProdCat}$ is bicategorically infinitely semi-additive (it has products, and it admits enrichment on the $2$-category of infinite monoidal categories, e.g. [35]).
Furthermore, we can extend the reasoning of Theorem 4.7 to the case of infinitary distributive and distributive categories; namely:
Theorem 4.8.
The free infinitary distributive category on the discrete category with two objects is equivalent to $\mathbf{Fam}\!\left(\mathbf{FinSet}^{\mathrm{op}}\times\mathbf{FinSet}^{\mathrm{op}}\right)$. Analogously, the free (finitary) distributive category on the discrete category with two objects is equivalent to $\mathbf{FinFam}\!\left(\mathbf{FinSet}^{\mathrm{op}}\times\mathbf{FinSet}^{\mathrm{op}}\right)$.